# Log-concave functions and transformations thereof

### v1

@@ Abstract
I summarize BagBerg05Logconcave’s review on log-concave functions, make several corrections, and augment the discussion with further results that can be useful in obtaining monotone hazard rate. I also provide an application of monopoly pricing.
@@ 3 Application: Log-Concave Demand Curve
There is a single product market. Consider a continuum of consumers, each of whom has one unit of indivisible demand of the product. Each consumer’s value of the unit of demand is $v\in[0,1]$ drawn i.i.d. from a distribution $G$ with strictly log-concave density gg. [^5] A weaker condition is to assume 1G(v)1-G(v) is strictly log-concave, with which all discussion in this section goes through.
@@
Remark. The fact that the price markup decreases in cost $c$ is a distinctive feature of log-concave demand. This property does not hold for general demand systems with concave revenue; for example, constant-elasticity demand exhibits increasing markup.
@@
Market power, elasticity, and information rent. Price markup is commonly used to measure market power: lower elasticity implies higher markup. Yet elasticity also mirrors the information environment. A less elastic (more inelastic) demand curve corresponds to a larger pool of high-value and relatively homogeneous consumers, which in turn implies a larger information rent for the indifferent consumer. Paradoxically, a higher information rent grants market power for the monopolist, even though in mechanism design it represents “difficulty” for the designer. The appropriate interpretation of this difficulty is thus the reluctance of supplying more good (and lowering price). Therefore, demand elasticity and information rent are two sides of the same coin. Both summarize the monopolist’s ability to extract surplus and thus provide dual interpretations of market power.

### v2

@@ Abstract
I summarize BagBerg05Logconcave’s review on log-concave functions, make several corrections, and augment the discussion with further results that can be useful in obtaining monotone hazard rate. I also provide an application of monopoly pricing, where strict log-concavity of demand curve implies strict concavity of the revenue function.
@@ 3 Application: Log-Concave Demand Curve
There is a single product market. Consider a continuum of consumers, each of whom has one unit of indivisible demand of the product. Each consumer’s value of the unit of demand is $v\in[0,1]$ drawn i.i.d. from a distribution $G$ with log-concave density gg. [^5] A weaker condition is to assume 1G(v)1-G(v) is strictly log-concave, with which all discussion in this section goes through.
@@
Remark. The fact that the price markup decreases in cost $c$ is a distinctive feature of log-concave demand. This property does not hold for general demand systems with concave revenue; for example, constant-elasticity demand (with elasticity larger than $1$) exhibits increasing markup.
@@
Market power, elasticity, and information rent. Price markup is commonly used to measure market power: lower elasticity implies higher markup. Yet elasticity also mirrors the information environment. A less elastic (more inelastic) demand curve corresponds to a larger pool of high-value and relatively homogeneous consumers, which in turn implies a larger information rent for the indifferent consumer. Paradoxically, a higher information rent grants market power for the monopolist, even though in mechanism design it represents “difficulty” for the designer. The appropriate interpretation of this difficulty is thus the reluctance of supplying more good (and lowering price) when making the marginal tradeoff. Therefore, demand elasticity and information rent are two sides of the same coin. Both summarize the monopolist’s ability to extract surplus and thus provide dual interpretations of market power.